# Isometric immersions into three-dimensional unimodular metric Lie groups

### v1

@@ Abstract.
We study isometric immersions of surfaces into simply connected $3$-dimensional unimodular Lie groups endowed with either Riemannian or Lorentzian left-invariant metrics, assuming that Milnor’s operator is diagonalizable in the Lorentzian case. We provide global models in coordinates for all these metric Lie groups that depend analytically on the structure constants and establish some fundamental theorems characterizing such immersions. In this sense, we study up to what extent we can recover the immersion from (a) the tangent projections of the natural left-invariant ambient frame, (b) the left-invariant Gauss map, and (c) the shape operator. As an application, we prove that an isometric immersion is determined by its left-invariant Gauss map up to certain well controlled angular companions. We also we classify totally geodesic surfaces and introduce Lorentzian analogues of the Daniel correspondence within two families of Lorentzian homogeneous $3$-manifolds with $4$-dimensional isometry group. We also classify isometric immersions in $\mathbb{R}^{3}$ or $\mathbb{S}^{3}$ whose left-invariant Gauss maps differ by a direct isometry of $\mathbb{S}^{2}$. Finally, we show that Daniel’s is the furthest extension of the classical Lawson correspondence within constant mean curvature surfaces in Riemannian unimodular metric Lie groups.
@@ 1. Introduction
It is a foundational problem in classical submanifold theory to decide if a given Riemannian manifold $M$ admits an isometric immersion $\phi:M\to\overline{M}$ into another Riemannian manifold $\overline{M}$ of higher dimension. The usual way of addressing this question is by finding geometric fundamental data that $M$ inherits from $\overline{M}$ via $\phi$, with the requirement that the existence of the immersion only depends on whether or not these data satisfy some differential equations on $M$, called fundamental equations. The earliest result of the kind dates back to Bonnet, who proved that the intrinsic metric and the shape operator (i.e., the first and second fundamental forms) of a surface immersed in Euclidean space $\mathbb{R}^{3}$ are enough to recover the immersion up to rigid motions, being Gauss and Codazzi equations the fundamental equations in this case. This was later generalized to immersions with arbitrary codimension in ambient spaces of constant curvature, namely $\mathbb{R}^{n}$, $\mathbb{S}^{n}$, and $\mathbb{H}^{n}$ (e.g., see [31, §7.C] and references therein).
@@
This family includes the Riemannian space forms $\mathbb{R}^{3}$ and $\mathbb{S}^{3}$, and in parallel the Minkowski space $\mathbb{L}^{3}$ and the universal cover of the Anti-de Sitter space $\mathbb{H}^{3}_{1}$. It also contains the Lorentzian version $\mathbb{L}(\kappa,\tau)$ of the family $\mathbb{E}(\kappa,\tau)$, see [17]. It is well known that both $\mathbb{E}(\kappa,\tau)$ and $\mathbb{L}(\kappa,\tau)$ are Killing submersions with constant bundle curvature $\tau$ over the complete Riemannian surface $\mathbb{M}^{2}(\kappa)$ of constant curvature $\kappa$, see also [18, 11]. Interestingly, among Lorentzian metric Lie groups of diagonalizable type, we can find another new family $\widehat{\mathbb{L}}(\kappa,\tau)$ of spaces that admit a semi-Riemannian submersion with constant bundle curvature $\tau$ over the complete Lorentzian surface $\mathbb{M}^{2}_{1}(\kappa)$ with constant curvature $\kappa$. Indeed, $\widehat{\mathbb{L}}(\kappa,\tau)$ also admits a unitary (spacelike) Killing vector field tangent to the fibers of the submersion.
@@
Acknowledgement. This work was supported by the research grant PID2022-142559NB-I00, funded by MCIN/AEI/10.13039/501100011033, and forms part of the third author’s PhD thesis. The authors thank M. Domínguez-Vázquez for his valuable comments, which improved the final version of the manuscript.
@@ 2. Unimodular metric Lie groups of diagonalizable type
[…]
$[X,Y]=L(X\times Y),\quad\text{for all }X,Y\in\mathfrak{g}.$ (2.1)
@@
It turns out that $G$ is unimodular if and only if $L$ in (2.1) is self-adjoint, see [28, Lem. 2.1]. We will further restrict to the case $L$ is diagonalizable (this is automatically true in the Riemannian case $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=1$), so we will say that $G$ is an unimodular metric Lie group of diagonalizable type. In other words, there exists a positively oriented left-invariant orthonormal frame $\{E_{1},E_{2},E_{3}\}$ in $G$ such that $\langle E_{i},E_{j}\rangle=\epsilon_{i}\delta_{j}^{i}$ and constants $c_{1},c_{2},c_{3}\in\mathbb{R}$ such that
@@
$+$ $+$ $+$ $\operatorname{SU}(2)$
$+$ $+$ -- $\widetilde{\mathrm{Sl}}_{2}(\mathbb{R})$
$+$ $+$ 00 $\widetilde{\mathrm{E}}(2)$
$+$ -- 00 $\mathrm{Sol}_{3}$
$+$ 00 00 $\mathrm{Nil}_{3}$
00 00 00 $\mathbb{R}^{3}$
@@
$+$ $+$ $+$ $\operatorname{SU}(2)$
$+$ $+$ -- $\widetilde{\mathrm{Sl}}_{2}(\mathbb{R})$
$+$ -- $+$ $\widetilde{\mathrm{Sl}}_{2}(\mathbb{R})$
$+$ $+$ 00 $\widetilde{\mathrm{E}}(2)$
$+$ 00 $+$ $\widetilde{\mathrm{E}}(2)$
$+$ -- 00 $\mathrm{Sol}_{3}$
$+$ 00 -- $\mathrm{Sol}_{3}$
$+$ 00 00 $\mathrm{Nil}_{3}$
00 00 $+$ $\mathrm{Nil}_{3}$
00 00 00 $\mathbb{R}^{3}$
@@ 2.1. The isometry group
¯$\begin{array}[]{lclcl}\overline{\nabla}_{E_{1}}E_{1}=0,&&\overline{\nabla}_{E_{1}}E_{2}=\epsilon_{3}\mu_{1}E_{3},&&\overline{\nabla}_{E_{1}}E_{3}=-\epsilon_{2}\mu_{1}E_{2},\\ \overline{\nabla}_{E_{2}}E_{1}=-\epsilon_{3}\mu_{2}E_{3},&&\overline{\nabla}_{E_{2}}E_{2}=0,&&\overline{\nabla}_{E_{2}}E_{3}=\epsilon_{1}\mu_{2}E_{1},\\ \overline{\nabla}_{E_{3}}E_{1}=\epsilon_{2}\mu_{3}E_{2},&&\overline{\nabla}_{E_{3}}E_{2}=-\epsilon_{1}\mu_{3}E_{1},&&\overline{\nabla}_{E_{3}}E_{3}=0.\end{array}$ (2.4)
@@
The following result follows by just checking that both sides coincide in the orthonormal frame $\{E_{1},E_{2},E_{3}\}$, which is a long but straightforward computation (see also [20, Lem. 2.2], where the Riemannian case is studied).
@@
Note that a metric Lie group has isometry group of dimension $3$, $4$ or $6$. We will now give a classification of the most symmetric cases.
@@
    If $\epsilon_{1}=1$ and $\epsilon_{2}=-1$, we can also set $c_{1}=-c_{2}=\frac{-\kappa}{2\tau}$ and $c_{3}=2\tau\neq 0$. This Lorentzian space $G$ admits a semi-Riemannian Killing submersion (in which the Killing vector field is spacelike) with constant bundle curvature $\tau$ over the complete simply connected Lorentzian surface $\mathbb{M}^{2}_{1}(\kappa)$ of constant curvature $\kappa$. We also get an isometry group of dimension $4$, since the $3$-dimensional group $\operatorname{Iso}(\mathbb{M}^{2}_{1}(\kappa))$ can be lifted to $G$ up to vertical translations with respect to the aforesaid submersion. Although the theory of semi-Riemannian Killing submersions is not established yet, these are standard arguments that can be straightforwardly adapted from [19, 18] to the Lorentzian case. Note that models given by Remark 2.4 enable explicit verifications of these assertions.
@@
    Assume now $\mu_{3}=0$ and $\mu_{1},\mu_{2}\neq 0$, which implies that $\lambda_{3}$ is not equal to $\lambda_{1}$ or $\lambda_{2}$, whence any isometry $\Phi\in\operatorname{Iso}(G)$ must preserve $E_{3}$ or send it to $-E_{3}$. Therefore, $\operatorname{Iso}(G)$ has dimension $4$ if and only if there is a continuous $1$-parameter subgroup in the stabilizer (which must preserve the orientation, the causality and the direction $E_{3}$). This directs our attention to isometries that rotate $E_{1}$ and $E_{2}$ (in a Riemannian or Lorentzian sense).
@@
The isometry group of a metric Lie group $G$ is determined by the left-translations by elements of $G$ itself and by the stabilizer $\operatorname{Stab}_{p}$ of any $p\in G$ (all stabilizers are conjugate by left-translations). It is well known that each element of $\operatorname{Stab}_{p}$ preserves left-invariant vector fields so it becomes an isometry of the Lie algebra $\mathfrak{g}$. Thus, $\operatorname{Stab}_{p}$ can be identified naturally with a subgroup of $\operatorname{O}(3)$ in the Riemannian case or $\operatorname{O}(2,1)$ in the Lorentzian case. If $\dim(\operatorname{Iso}(G))=6$, then $\operatorname{Stab}_{p}$ is well known is all $\operatorname{O}(3)$ or $\operatorname{O}(2,1)$. If $\dim(\operatorname{Iso}(G))=4$, then $\operatorname{Stab}_{p}$ contains a Lie subgroup $\operatorname{SO}(2)$ or $\operatorname{SO}(1,1)$ of rotations about the unit Killing direction. If $\dim(\operatorname{Iso}(G))=3$, then $\operatorname{Stab}_{p}$ is discrete and contains either $4$ or $8$ elements, depending on whether or not $G$ admits orientation-reversing isometries (see [23, Props. 2.21 and 2.24], although the proof easily extends to the semi-Riemannian case since elements of $\operatorname{Stab}_{p}$ must preserve the directions that diagonalize the Ricci operator by the same arguments as in the proof of Proposition 2.2).
@@ 2.2. A unified model in coordinates
where $\lambda=\bigl{(}1+\tfrac{\kappa}{4}(x^{2}+y^{2})\bigr{)}^{-1}$. The sign of $\tau$ is chosen to agree with the definition of bundle curvature with the standard orientation and unit Killing $\partial_{z}$ (cf. [11, Ex. 1.4]).
@@
where $\lambda=(1+\frac{\kappa}{4}(x^{2}-y^{2}))^{-1}$. Notice that $E_{1}=\tfrac{1}{\lambda}\partial_{x}-\tau y\partial_{z}$, $E_{2}=\tfrac{1}{\lambda}\partial_{y}+\tau x\partial_{z}$ and $E_{3}=\partial_{z}$ form a global orthonormal frame with respect to (2.12), where $E_{2}$ is timelike and $E_{3}$ is a spacelike unit Killing vector field. The natural submersion $(x,y,z)\mapsto(x,y)$ is semi-Riemannian with constant bundle curvature $\tau=\frac{1}{2}\langle[E_{1},E_{2}],E_{3}\rangle$ (cf. [11, Eqn. (2.10)]) over the base surface $D$ equipped with the Lorentzian metric $\lambda^{2}(\mathrm{d}x^{2}-\mathrm{d}y^{2})$ of constant curvature $\kappa$. This calculation of the curvature follows from [25, Prop. 3.44], for we can work out
@@
Our approach to these models is in turn based on the canonical frame for Killing submersions, which uses $\partial_{z}$ as the distinguished Killing vector field, see [19, 18, 11]. The algebraic Killing vector fields in a metric Lie group are the right-invariant vector fields; in constrast, our model for metric Lie groups takes $\partial_{z}$ as a left-invariant vector field that diagonalizes the Ricci operator. Our approach has the advantage that even in the same Lie group we have three models depending on which of the directions $E_{1},E_{2},E_{3}$ is chosen to become $\partial_{z}$, and this amounts to rotate the indexes cyclically. The rest of this subsection will be devoted to analyze if the models are global. To this end, we will begin by discussing the relation between our models and the traditional models for the three big families of unimodular Lie groups (namely, $\widetilde{\mathrm{SL}}_{2}(\mathbb{R})$, $\mathrm{SU}(2)$ and $\mathbb{R}^{2}\rtimes_{A}\mathbb{R}$).
@@
If $c_{1},c_{2}>0$ and $c_{3}<0$, then $G$ is $\widetilde{\mathrm{SL}}_{2}(\mathbb{R})$, the universal cover of $\mathrm{SL}_{2}(\mathbb{R})$, equipped with a left-invariant metric. The group $\mathrm{SL}_{2}(\mathbb{R})$ is usually represented by the matrices $(\begin{smallmatrix}a&b\\ \bar{b}&\bar{a}\end{smallmatrix})\in\mathcal{M}_{2}(\mathbb{C})$ with determinant $1$ (the product in $\mathrm{SL}_{2}(\mathbb{R})$ is just the product of matrices), so that it can be identified with the quadric $Q\equiv\{(a,b)\in\mathbb{C}^{2}:|a|^{2}-|b|^{2}=1\}$ that represents the unit sphere in $\mathbb{R}^{4}_{2}$. The frame given by $X_{1}=\frac{1}{2}\sqrt{c_{2}c_{3}}(i\bar{b},i\bar{a})$, $X_{2}=\frac{1}{2}\sqrt{c_{1}c_{3}}(\bar{b},\bar{a})$ and $X_{3}=\frac{1}{2}\sqrt{c_{2}c_{3}}(ia,ib)$ is left-invariant with $[X_{1},X_{2}]=c_{3}X_{3}$, $[X_{2},X_{3}]=c_{1}X_{1}$ and $[X_{3},X_{1}]=c_{2}X_{2}$ and the map $F:D\times\mathbb{R}\to Q$ given by
@@
If $c_{1},c_{2},c_{3}>0$, then $G$ is isomorphic and isometric to $\operatorname{SU}(2)$ with a left-invariant metric. This group contains (by definition) the unitary matrices $(\begin{smallmatrix}a&b\\ -\bar{b}&\bar{a}\end{smallmatrix})\in\mathcal{M}_{2}(\mathbb{C})$ with determinant $1$, when the product of matrices is considered. This yields an identification with $\mathbb{S}^{3}\equiv\{(a,b)\in\mathbb{C}^{2}:|a|^{2}+|b|^{2}=1\}$, where the vector fields $X_{1}=\frac{1}{2}\sqrt{c_{2}c_{3}}(-i\bar{b},i\bar{a})$, $X_{2}=\frac{1}{2}\sqrt{c_{1}c_{3}}(-\bar{b},\bar{a})$ and $X_{3}=\frac{1}{2}\sqrt{c_{2}c_{3}}(ia,ib)$ become left-invariant with $[X_{1},X_{2}]=c_{3}X_{3}$, $[X_{2},X_{3}]=c_{1}X_{1}$ and $[X_{3},X_{1}]=c_{2}X_{2}$. The map $F:D\times\mathbb{R}\to\mathbb{S}^{3}-\{(e^{i\theta},0):\theta\in\mathbb{R}\}$ given by
@@
In view of [23, §2.2], this means that $X_{1}=c(z)\partial_{x}+c_{2}s(z)\partial_{y}$, $X_{2}=-c_{1}s(z)\partial_{x}+c(z)\partial_{y}$ and $X_{3}=\partial_{z}$ define a global left-invariant frame for $G$, viewed as $\mathbb{R}^{3}\equiv\mathbb{R}^{2}\rtimes_{A}\mathbb{R}$. Since $c_{3}=0$, we get from (2.9) that $X_{i}=E_{i}$ for all $i$, whence the metric (2.10) coincides with the standard metric as a semidirect product.

### v2

@@ Abstract.
We study isometric immersions of surfaces into simply connected $3$-dimensional unimodular Lie groups endowed with either Riemannian or Lorentzian left-invariant metrics, assuming that Milnor’s operator is diagonalizable in the Lorentzian case. We provide global models in coordinates for all these metric Lie groups that depend analytically on the structure constants and establish some fundamental theorems characterizing such immersions. In this sense, we study up to what extent we can recover the immersion from (a) the tangent projections of the natural left-invariant ambient frame, (b) the left-invariant Gauss map, and (c) the shape operator. In particular, we show that an isometric immersion is determined by its left-invariant Gauss map up to certain well controlled angular companions. We also classify totally geodesic surfaces and introduce four Lorentzian analogues of the Daniel correspondence within two families of Lorentzian homogeneous $3$-manifolds with $4$-dimensional isometry group. We also classify isometric immersions in $\mathbb{R}^{3}$ or $\mathbb{S}^{3}$ whose left-invariant Gauss maps differ by a direct isometry of $\mathbb{S}^{2}$. Finally, we show that Daniel’s is the furthest extension of the classical Lawson correspondence for constant mean curvature surfaces in Riemannian unimodular metric Lie groups.
@@ 1. Introduction
It is a foundational problem in classical submanifold theory to decide if a given Riemannian manifold $M$ admits an isometric immersion $\phi:M\to\overline{M}$ into another Riemannian manifold $\overline{M}$ of higher dimension. The usual way of addressing this question is by finding geometric fundamental data that $M$ inherits from $\overline{M}$ via $\phi$, with the requirement that the existence of the immersion only depends on whether or not these data satisfy some differential equations on $M$, called fundamental equations. The earliest result of the kind dates back to Bonnet, who proved that the intrinsic metric and the shape operator (i.e., the first and second fundamental forms) of a surface immersed in Euclidean space $\mathbb{R}^{3}$ are enough to recover the immersion up to rigid motions, being Gauss and Codazzi equations the fundamental equations in this case. This was later generalized to immersions with arbitrary codimension in ambient spaces of constant curvature, namely $\mathbb{R}^{n}$, $\mathbb{S}^{n}$, and $\mathbb{H}^{n}$ (e.g., see [32, §7.C] and references therein).
@@
This family includes the Riemannian space forms $\mathbb{R}^{3}$ and $\mathbb{S}^{3}$, and in parallel the Minkowski space $\mathbb{L}^{3}$ and the universal cover of the Anti-de Sitter space $\mathbb{H}^{3}_{1}$. It also contains the Lorentzian version $\mathbb{L}(\kappa,\tau)$ of the family $\mathbb{E}(\kappa,\tau)$, see [18]. It is well known that both $\mathbb{E}(\kappa,\tau)$ and $\mathbb{L}(\kappa,\tau)$ are unit Killing submersions with constant bundle curvature $\tau$ over the complete Riemannian surface $\mathbb{M}^{2}(\kappa)$ of constant curvature $\kappa$, see also [19, 9]. Interestingly, among Lorentzian metric Lie groups of diagonalizable type, we can find another new family $\widehat{\mathbb{L}}(\kappa,\tau)$ of spaces that admit a semi-Riemannian submersion with constant bundle curvature $\tau$ over the complete Lorentzian surface $\mathbb{M}^{2}_{1}(\kappa)$ with constant curvature $\kappa$. Indeed, $\widehat{\mathbb{L}}(\kappa,\tau)$ also admits a unitary (spacelike) Killing vector field tangent to the fibers of the submersion. After the submission of this work, we found this new family of spaces was already recently discovered independently by Calvaruso and Pellegrino [4]. Note also that there are other Lorentzian metric Lie groups with isometry group of dimension 4, see [4, Rmk. 2.5].
@@
Acknowledgement. This work was supported by the research grant PID2022-142559NB-I00, funded by MCIN/AEI/10.13039/501100011033, and forms part of the third author’s PhD thesis. The authors would like to thank B. Daniel and M. Domínguez-Vázquez for their valuable comments, which improved the final version of the manuscript.
@@ 2. Unimodular metric Lie groups of diagonalizable type
A Lie group $G$ is called unimodular if its left-invariant Haar measure is also right-invariant. It is well known that $G$ is unimodular if, and only if, for any $X\in\mathfrak{g}$, the adjoint endomorphism $\mathrm{ad}_{X}:\mathfrak{g}\to\mathfrak{g}$ given by $\mathrm{ad}_{X}(Y)=[X,Y]$ has trace equal to zero (as usual, $\mathfrak{g}$ denotes the Lie algebra of $G$). If $G$ has dimension $3$, it is oriented and endowed with a (nondegenerate) semi-Riemannian metric $\langle\cdot,\cdot\rangle$, both the cross product $\times$ and the Lie bracket $\cdot$$\cdot$] are skew-symmetric, Milnor [25] considered the unique linear operator $L:\mathfrak{g}\to\mathfrak{g}$ such that
$X,Y]=L(X\times Y),\quad\text{for all }X,Y\in\mathfrak{g}.$ (2.1)
@@
It turns out that $G$ is unimodular if and only if $L$ in (2.1) is self-adjoint, see [29, Lem. 2.1]. We will further restrict to the case $L$ is diagonalizable (this is automatically true in the Riemannian case $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=1$), so we will say that $G$ is an unimodular metric Lie group of diagonalizable type. In other words, there exists a positively oriented left-invariant orthonormal frame $\{E_{1},E_{2},E_{3}\}$ in $G$ such that $\langle E_{i},E_{j}\rangle=\epsilon_{i}\delta_{j}^{i}$ and constants $c_{1},c_{2},c_{3}\in\mathbb{R}$ such that
@@
++ ++ ++ $\operatorname{SU}(2)$
++ ++ - $\widetilde{\mathrm{Sl}}_{2}(\mathbb{R})$
++ ++ 0 $\widetilde{\mathrm{E}}(2)$
++ - 0 $\mathrm{Sol}_{3}$
++ 0 0 $\mathrm{Nil}_{3}$
0 0 0 $\mathbb{R}^{3}$
@@
++ ++ ++ $\operatorname{SU}(2)$
++ ++ - $\widetilde{\mathrm{Sl}}_{2}(\mathbb{R})$
++ - ++ $\widetilde{\mathrm{Sl}}_{2}(\mathbb{R})$
++ ++ 0 $\widetilde{\mathrm{E}}(2)$
++ 0 ++ $\widetilde{\mathrm{E}}(2)$
++ - 0 $\mathrm{Sol}_{3}$
++ 0 - $\mathrm{Sol}_{3}$
++ 0 0 $\mathrm{Nil}_{3}$
0 0 ++ $\mathrm{Nil}_{3}$
0 0 0 $\mathbb{R}^{3}$
@@ 2.1. The isometry group
$\begin{array}[]{lclcl}\overline{\nabla}_{E_{1}}E_{1}=0,&&\overline{\nabla}_{E_{1}}E_{2}=\epsilon_{3}\mu_{1}E_{3},&&\overline{\nabla}_{E_{1}}E_{3}=-\epsilon_{2}\mu_{1}E_{2},\\ \overline{\nabla}_{E_{2}}E_{1}=-\epsilon_{3}\mu_{2}E_{3},&&\overline{\nabla}_{E_{2}}E_{2}=0,&&\overline{\nabla}_{E_{2}}E_{3}=\epsilon_{1}\mu_{2}E_{1},\\ \overline{\nabla}_{E_{3}}E_{1}=\epsilon_{2}\mu_{3}E_{2},&&\overline{\nabla}_{E_{3}}E_{2}=-\epsilon_{1}\mu_{3}E_{1},&&\overline{\nabla}_{E_{3}}E_{3}=0.\end{array}$ (2.4)
@@
The following result follows by just checking that both sides coincide in the orthonormal frame $\{E_{1},E_{2},E_{3}\}$, which is a long but straightforward computation (see also [21, Lem. 2.2], where the Riemannian case is studied).
@@
Note that a metric Lie group has isometry group of dimension $3$, 44 or $6$. We will now give a classification of the most symmetric cases.
@@
    If $\epsilon_{1}=1$ and $\epsilon_{2}=-1$, we can also set $c_{1}=-c_{2}=\frac{-\kappa}{2\tau}$ and $c_{3}=2\tau\neq 0$. This Lorentzian space $G$ admits a semi-Riemannian Killing submersion (in which the Killing vector field is spacelike) with constant bundle curvature $\tau$ over the complete simply connected Lorentzian surface $\mathbb{M}^{2}_{1}(\kappa)$ of constant curvature $\kappa$. We also get an isometry group of dimension $4$, since the $3$-dimensional group $\operatorname{Iso}(\mathbb{M}^{2}_{1}(\kappa))$ can be lifted to $G$ up to vertical translations with respect to the aforesaid submersion. Although the theory of semi-Riemannian Killing submersions is not established yet, these are standard arguments that can be straightforwardly adapted from [19] to the Lorentzian case. Note that models given by Remark 2.4 enable explicit verifications of these assertions.
@@
    Assume now $\mu_{3}=0$ and $\mu_{1},\mu_{2}\neq 0$, which implies that $\lambda_{3}$ is not equal to $\lambda_{1}$ or $\lambda_{2}$, whence any isometry $\Phi\in\operatorname{Iso}(G)$ must preserve $E_{3}$ or send it to $E_{3}$. Therefore, $\operatorname{Iso}(G)$ has dimension $4$ if and only if there is a continuous $1$-parameter subgroup in the stabilizer (which must preserve the orientation, the causality and the direction $E_{3}$). This directs our attention to isometries that rotate $E_{1}$ and $E_{2}$ (in a Riemannian or Lorentzian sense).
@@
The isometry group of a metric Lie group $G$ is determined by the left-translations by elements of $G$ itself and by the stabilizer $\operatorname{Stab}_{p}$ of any $p\in G$ (all stabilizers are conjugate by left-translations). It is well known that each element of $\operatorname{Stab}_{p}$ preserves left-invariant vector fields so it becomes an isometry of the Lie algebra $\mathfrak{g}$. Thus, $\operatorname{Stab}_{p}$ can be identified naturally with a subgroup of $\operatorname{O}(3)$ in the Riemannian case or $\operatorname{O}(2,1)$ in the Lorentzian case. If $\dim(\operatorname{Iso}(G))=6$, then $\operatorname{Stab}_{p}$ is well known is all $\operatorname{O}(3)$ or $\operatorname{O}(2,1)$. If $\dim(\operatorname{Iso}(G))=4$, then $\operatorname{Stab}_{p}$ contains a Lie subgroup $\operatorname{SO}(2)$ or $\operatorname{SO}(1,1)$ of rotations about the unit Killing direction. If $\dim(\operatorname{Iso}(G))=3$, then $\operatorname{Stab}_{p}$ is discrete and contains either $4$ or $8$ elements, depending on whether or not $G$ admits orientation-reversing isometries (see [24, Props. 2.21 and 2.24], although the proof easily extends to the semi-Riemannian case since elements of $\operatorname{Stab}_{p}$ must preserve the directions that diagonalize the Ricci operator by the same arguments as in the proof of Proposition 2.2).
@@ 2.2. A unified model in coordinates
where $\lambda=\bigl(1+\tfrac{\kappa}{4}(x^{2}+y^{2})\bigr)^{-1}$. The sign of $\tau$ is chosen to agree with the definition of bundle curvature with the standard orientation and unit Killing $\partial_{z}$ (cf. [9, Ex. 1.4]).
@@
where $\lambda=(1+\frac{\kappa}{4}(x^{2}-y^{2}))^{-1}$. Notice that $E_{1}=\tfrac{1}{\lambda}\partial_{x}-\tau y\partial_{z}$, $E_{2}=\tfrac{1}{\lambda}\partial_{y}+\tau x\partial_{z}$ and $E_{3}=\partial_{z}$ form a global orthonormal frame with respect to (2.12), where $E_{2}$ is timelike and $E_{3}$ is a spacelike unit Killing vector field. The natural submersion $(x,y,z)\mapsto(x,y)$ is semi-Riemannian with constant bundle curvature $\tau=\frac{1}{2}\langle[E_{1},E_{2}],E_{3}\rangle$ (cf. [9, Eqn. (2.10)]) over the base surface $D$ equipped with the Lorentzian metric $\lambda^{2}(\mathrm{d}x^{2}-\mathrm{d}y^{2})$ of constant curvature $\kappa$. This calculation of the curvature follows from [26, Prop. 3.44], for we can work out
@@
Our approach to these models is in turn based on the canonical frame for Killing submersions, which uses $\partial_{z}$ as the distinguished Killing vector field, see [19, 9]. The algebraic Killing vector fields in a metric Lie group are the right-invariant vector fields; in constrast, our model for metric Lie groups takes $\partial_{z}$ as a left-invariant vector field that diagonalizes the Ricci operator. Our approach has the advantage that even in the same Lie group we have three models depending on which of the directions $E_{1},E_{2},E_{3}$ is chosen to become $\partial_{z}$, and this amounts to rotate the indexes cyclically. The rest of this subsection will be devoted to analyze if the models are global. To this end, we will begin by discussing the relation between our models and the traditional models for the three big families of unimodular Lie groups (namely, $\widetilde{\mathrm{SL}}_{2}(\mathbb{R})$, $\mathrm{SU}(2)$ and $\mathbb{R}^{2}\rtimes_{A}\mathbb{R}$).
@@
If $c_{1},c_{2}>0$ and $c_{3}<0$, then $G$ is $\widetilde{\mathrm{SL}}_{2}(\mathbb{R})$, the universal cover of $\mathrm{SL}_{2}(\mathbb{R})$, equipped with a left-invariant metric. The group $\mathrm{SL}_{2}(\mathbb{R})$ is usually represented by the matrices $(\begin{smallmatrix}a&b\\ \bar{b}&\bar{a}\end{smallmatrix})\in\mathcal{M}_{2}(\mathbb{C})$ with determinant $1$ (the product in $\mathrm{SL}_{2}(\mathbb{R})$ is just the product of matrices), so that it can be identified with the quadric $Q\equiv\{(a,b)\in\mathbb{C}^{2}:|a|^{2}-|b|^{2}=1\}$ that represents the unit sphere in $\mathbb{R}^{4}_{2}$. The frame given by $X_{1}=\frac{1}{2}\sqrt{c_{2}c_{3}}(i\bar{b},i\bar{a})$, $X_{2}=\frac{1}{2}\sqrt{c_{1}c_{3}}(\bar{b},\bar{a})$ and $X_{3}=\frac{1}{2}\sqrt{c_{2}c_{3}}(ia,ib)$ is left-invariant with $X_{1},X_{2}]=c_{3}X_{3}$, $X_{2},X_{3}]=c_{1}X_{1}$ and $X_{3},X_{1}]=c_{2}X_{2}$ and the map $F:D\times\mathbb{R}\to Q$ given by
@@
If $c_{1},c_{2},c_{3}>0$, then $G$ is isomorphic and isometric to $\operatorname{SU}(2)$ with a left-invariant metric. This group contains (by definition) the unitary matrices $(\begin{smallmatrix}a&b\\ -\bar{b}&\bar{a}\end{smallmatrix})\in\mathcal{M}_{2}(\mathbb{C})$ with determinant $1$, when the product of matrices is considered. This yields an identification with $\mathbb{S}^{3}\equiv\{(a,b)\in\mathbb{C}^{2}:|a|^{2}+|b|^{2}=1\}$, where the vector fields $X_{1}=\frac{1}{2}\sqrt{c_{2}c_{3}}(-i\bar{b},i\bar{a})$, $X_{2}=\frac{1}{2}\sqrt{c_{1}c_{3}}(-\bar{b},\bar{a})$ and $X_{3}=\frac{1}{2}\sqrt{c_{2}c_{3}}(ia,ib)$ become left-invariant with $X_{1},X_{2}]=c_{3}X_{3}$, $X_{2},X_{3}]=c_{1}X_{1}$ and $X_{3},X_{1}]=c_{2}X_{2}$. The map $F:D\times\mathbb{R}\to\mathbb{S}^{3}-\{(e^{i\theta},0):\theta\in\mathbb{R}\}$ given by
@@
In view of [24, §2.2], this means that $X_{1}=c(z)\partial_{x}+c_{2}s(z)\partial_{y}$, $X_{2}=-c_{1}s(z)\partial_{x}+c(z)\partial_{y}$ and $X_{3}=\partial_{z}$ define a global left-invariant frame for $G$, viewed as $\mathbb{R}^{3}\equiv\mathbb{R}^{2}\rtimes_{A}\mathbb{R}$. Since $c_{3}=0$, we get from (2.9) that $X_{i}=E_{i}$ for all $i$, whence the metric (2.10) coincides with the standard metric as a semidirect product.